OpenStudy (anonymous):

I need help solving this system: 7x^2-9y^2=31 and 5x^2+9y^2=161. It would be nice to have an explanation too.

OpenStudy (anonymous):

add them as a start. the y^2 will drop out and you will have a quadratic in x

OpenStudy (anonymous):

ok so we have: \[7x ^{2}-9y ^{2}=31\] \[5x ^{2}+9y ^{2}=161\] add and we get:\[12x ^{2}=192\] divide by 12 \[x^2=16\] \[x=4\] plug in to find y \[7(4)^2-9y^2=31\] \[112-9y^2=31\] \[-9y^2=-81\] \[y^2= 9\] \[y=3\]